OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6

OpenStudy (lgbasallote):

use the two point formula \[\large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x - x_1)\] \[\implies y - 2 = \frac{3 - 2}{6-2} (x - 2)\] \[\implies y- 2 = \frac 14 (x-2)\] multiply all terms by 4 \[\implies 4y - 8 = x - 2\] put everything to the left side \[\implies 4y - 8 - x + 2 = 0\] \[\implies 4y - x - 6 = 0\] does that help?

OpenStudy (lgbasallote):

btw if you think i just gave you an answer weely-neely you're wrong...you have to manipulate that thing i gave so it'll look something like one of your choices ;)